# Complexity-aware fine-tuning

### v1

@@ 4.3 Complexity estimation evaluation
4.3.1 Single token and chain-of-thought answer entropy
Tables 2 and 3 present ROC AUC and accuracy for single token entropy and for the entropy of the answer token in the chain-of-thought type of response, respectively. Metrics are calculated for the categories provided by MMLU-Pro as well as education levels and reasoning scores estimated via MASJ.
IDK responses and results with invalid formatting are excluded from the calculations.
We observe that the accuracy tends to be slightly higher when an LLM can answer ’IDK’. At the same time, ’IDK’ responses do not consistently affect ROC AUC for all models.
@@ 4.3.1 Single token and chain-of-thought answer entropy
4.3.2 Chain-of-thought aggregated response entropy
Method Qwen 3B Qwen 3B* Phi4-mini Phi4-mini*
Answer Entropy (AE) 0.68 0.67 0.61 0.58
COT Mean 0.59 0.58 0.59 0.63
COT Max 0.63 0.61 0.6 0.65
Sequence Mean Mean 0.6 0.59 0.6 0.62
Sequence Max Mean 0.59 0.58 0.59 0.61
Sequence Mean Max 0.62 0.6 0.59 0.62
Marginal Diff Mean 0.58 0.57 0.58 0.61
Top-2 Entropies Diff 0.51 0.5 0.5 0.51
COT Max minus AE 0.54 0.53 0.51 0.57
COT Max and AE 0.7 0.69 0.62 0.62
Number of Samples 11049 10724 9997 9973
Table 4: ROC AUC values for COT response
* Alternative prompt to allow model answer "I do not know"
Table 4 provides results, which highlight our key observations:
  • Simple answer entropy (token with the chosen option ID at the end of a chain-of-thought) often outperforms more complex COT aggregation methods, particularly in models with strong baseline performance (e.g., Qwen-3B achieving $0.68$ ROC AUC). Although the hybrid method outperforms the answer entropy, the main weights of the hybrid linear combination were assigned to the answer entropy.
5 Conclusion and discussion
Using the entropy-based data split, we find that different complexity scores require different training approaches. Standard supervised fine-tuning (SFT) is enough for the easy and medium bands, but lags on the hard band. For the hard questions, adding a distilled chain-of-thought from a large LLM unlocks further gains (accuracies of 0.5/0.6 vs. 0.4/0.46 for Qwen 3B/Phi-4-mini), while providing performance comparable with training on the distilled dataset without the data split while using $62\%$ less data (accuracies of 0.5/0.6 vs. 0.49/0.62 for Qwen 3B/Phi-4-mini).
We publish collected data and code to support further research in the area of numerical complexity estimation.
Limitations
  • In low-resource settings teacher may ßbe unavailable or imperfect, which reduces the benefit of learning from a distilled chain-of-thought. Additionally, we did not explore how well the approach generalizes to other reasoning-promoting techniques.
  • We did not run the pipeline for more epochs due to resource limitations, so its behavior for the longer runs stays unknown.
Appendix A Prompts
A.1 Prompts used for complexity estimation
System prompt for a single token response
System prompt for a chain-of-thought response
The following are multiple choice questions about subject. Explain your thinking process step-by-step. At the end, write down the number of the correct answer by strictly following this format: $[[\text{number of correct answer}]]$.
The following are multiple choice questions about subject. Explain your thinking process step-by-step. At the end, if you are certain about the answer write down the number of the correct answer by strictly following this format: $[[\text{number of correct answer}]]$, otherwise return $[[\text{0}]]$.
The following are multiple choice questions about subject. Explain your thinking process step-by-step. At the end, if you know the answer write down the number of the correct answer by strictly following this format: $[[\text{number of correct answers}]]$, otherwise return $[[\text{0}]]$.
User prompt
Question: …
Options:
1. …
2. …
n. …
Table 5: Prompts for complexity estimation
A.2 General prompts
Table 6: Prompt for MASJ reasoning
Table 7: Prompt for MASJ education levels
Appendix B Aggregation Methods
Our analysis considers the following methods:
  1. 1.
    CoT word-aggregation methods
    • Single Token Answer Entropy
    • CoT Mean
    • CoT Max
    • Difference between CoT Max and Single Token Answer Entropy
  2. 2.
    CoT sequence-aggregation methods
    • Sequence Mean of Words Mean
    • Sequence Max of Words Mean
    • Sequence Mean of Words Max
  3. 3.
    Probability-based methods
  4. 4.
    Hybrid method
B.1 Word-aggregation Methods
This CoT aggregations have the same entropy values as in 3.2.3 for each CoT token.
$h_{j}=-\sum_{i=1}^{n}p_{i}\log{p_{i}}$
where $p_{i}$ probability of a single token, $n$ vocabulary size, $h_{j}$ entropy of the corresponded token, $h_{answer}$ entropy of the answer token, and $N$ is the token amount in LLM response.
$CoT_{mean}=\sum_{j=1}^{N}\frac{h_{j}}{N}$
$CoT_{max}=\max_{j}{h_{j}}$
So, Chain-of-Thought maximum and answer entropy difference is:
$|CoT_{max}-h_{answer}|$
B.2 Sequence-aggregation Methods
For $M$ logical claims, which were split by tokens that corresponded to the end of the sequence, we have tokens sets ${C_{1},C_{2}\ ...\ C_{M}}$.
$Seq_{mean}=\frac{1}{M}\sum_{j=1}^{M}[\frac{1}{|C_{j}|}\sum_{i\in C_{j}}h_{i}]$
$Seq_{mean,max}=\frac{1}{M}\sum_{j=1}^{M}[\max_{i\in C_{j}}h_{i}]$
$Seq_{max,mean}=\max_{j}[\frac{1}{|C_{j}|}\sum_{i\in C_{j}}h_{i}]$
B.3 Probability-based Methods
$\sigma_{i}=p_{i}^{1}-p_{i}^{2}$
$\overline{\sigma}=\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}$
Top-2 entropies difference in response $\delta$.
$\delta=\max_{j}h_{j}-\max_{i|i\neq j}h_{i}$
B.4 Hybrid Method
In this section we provide linear combination of $2$ best-perform previous methods: $h_{answer}$ and $CoT_{max}$. Also, we tried adding the third element $CoT_{mean}$, but it has decreased the ROC-AUC, so we made a decision to remove it.
$h_{mix}=(1-\alpha)*h_{answer}+\alpha*CoT_{max}$
where $0\leq\alpha\leq 1$ is the hyperparameter (e.g. $\alpha=0.05$ empirically for Qwen-3B).
@@ C.2 Feature importances of reasoning model
We evaluated logistic regression weights, that reflect feature importance. Our classifier achieves 0.721, 0.717 and 0.731 accuracies by using thinking total entropy, length of the reasoning chain or both features combined.

### v2

@@ 4.3.3 Quality of complexity estimation
Single token and chain-of-thought answer entropy
Table 3 presents the main ROC AUC values for single token entropy response and for the various aggregates of the chain-of-thought type of response. IDK responses and results with invalid formatting are excluded from the calculations.
We see there that the single token response shows the best results. In-depth analysis of ROC AUC and accuracy scores with splits across domains and larger selection of models can be found in tables 7 and 8 in the Appendix.
Key observations:
  • Single token response provides the best ROC AUC score. At the same time, ’IDK’ responses do not consistently affect ROC AUC for all models.
  • Accuracy tends to be slightly higher when an LLM can answer ’IDK’.
@@
5 Conclusion and discussion
Using the entropy-based data split, we find that different complexity scores require different training approaches. Standard supervised fine-tuning (SFT) is enough for the easy and medium bands, but lags on the hard band. For the hard questions, adding a distilled chain-of-thought from a large LLM unlocks further gains. Our pipeline achieves accuracies of 0.52/0.64 vs. 0.39/0.51 for Qwen 3B/Phi-4-mini, while using $81\%$ less data.
Limitations
  • In low-resource settings teacher may be unavailable or imperfect, which reduces the benefit of learning from a distilled chain-of-thought. Additionally, we did not explore how well the approach generalizes to other reasoning-promoting techniques.
Appendix A Prompts
A.1 Prompts used for complexity estimation
System prompt for a single token response
System prompt for a chain-of-thought response
The following are multiple choice questions about subject. Explain your thinking process step-by-step. At the end, write down the number of the correct answer by strictly following this format: [[number of correct answer]]$[$$\text${number of correct answer}]].
The following are multiple choice questions about subject. Explain your thinking process step-by-step. At the end, if you are certain about the answer write down the number of the correct answer by strictly following this format: [[number of correct answer]]$[$$\text${number of correct answer}]], otherwise return $\text{0}]]$.
The following are multiple choice questions about subject. Explain your thinking process step-by-step. At the end, if you know the answer write down the number of the correct answer by strictly following this format: [[number of correct answers]]$[$$\text${number of correct answers}]], otherwise return $\text{0}]]$.
User prompt
Question: …
Options:
1. …
2. …
n. …
Table 4: Prompts for complexity estimation
A.2 General prompts
Table 5: Prompt for MASJ reasoning
Table 6: Prompt for MASJ education levels
Appendix B Aggregation Methods
Our analysis considers diverse methods to identify the complexity of a questions based on outputs logits with a specific focus on commonly used entropy. They are listed in the list with more details provided below:
  1. 1.
    CoT word-aggregation methods
    • Single Token Answer Entropy
    • CoT Mean
    • CoT Max
    • Difference between CoT Max and Single Token Answer Entropy
  2. 2.
    CoT sequence-aggregation methods
    • Sequence Mean of Words Mean
    • Sequence Max of Words Mean
    • Sequence Mean of Words Max
  3. 3.
    Probability-based methods
  4. 4.
    Hybrid method
B.1 Word-aggregation Methods
This CoT aggregations have the same entropy values as in Section 3.2.3 for each CoT token:
$h_{j}=-\sum_{i=1}^{n}p_{i}\log{p_{i}},$
where $p_{i}$ is the probability of a specific token, $n$ is the vocabulary size, and $h_{j}$ is the entropy of the corresponded token. Aggregating per-token entropies for an answer of length $N$ get:
$\mathrm{CoT}_{\mathrm{mean}}$ $=\frac{1}{N}\sum_{j=1}^{N}h_{j},$
$\mathrm{CoT}_{\mathrm{max}}$ $=\max_{j}h_{j}.$
So, Chain-of-Thought maximum and answer entropy difference is:
$|\mathrm{CoT}_{\mathrm{max}}-h_{\mathrm{answer}}|,$
where $h_{\mathrm{answer}}$ is the entropy of the answer token.
B.2 Sequence-aggregation Methods
For $M$ logical claims, which were split by tokens that corresponded to the end of the sequence, we have tokens sets $C_{1},C_{2},\ldots,C_{M}$.
$\mathrm{Seq}_{\mathrm{mean}}$ $=\frac{1}{M}\sum_{j=1}^{M}\left[\frac{1}{|C_{j}|}\sum_{i\in C_{j}}h_{i}\right],$
$\mathrm{Seq}_{\mathrm{mean,max}}$ $=\frac{1}{M}\sum_{j=1}^{M}\left[\max_{i\in C_{j}}h_{i}\right],$
$\mathrm{Seq}_{\mathrm{max,mean}}$ $=\max_{j}\left[\frac{1}{|C_{j}|}\sum_{i\in C_{j}}h_{i}\right].$
B.3 Probability-based Methods
$\sigma_{i}$ $=p_{i}^{(}1)-p_{i}^{(}2),$
$\overline{\sigma}$ $=\frac{1}{N}\sum_{i=1}^{N}\sigma_{i},$
here $p_{i}^{(}k)$ is the $k$-th top value in the vector $\mathbf{p}_{i}$ of the probability distribution of the tokens from the dictionary during the generation of $i$-th token.
We can also consider differences between top two entropies in response $\delta$:
$\delta=\max_{j}h_{j}-\max_{i|i\neq j}h_{i}.$
B.4 Hybrid Method
We can also use a linear combination of $2$ best-perform previous methods: $h_{\mathrm{answer}}$ and $\mathrm{CoT}_{\max}$. Also, we tried adding the third element $\mathrm{CoT}_{\mathrm{mean}}$, but it has decreased the ROC-AUC, so we made a decision to remove it.
$h_{\mathrm{mix}}=(1-\alpha)h_{\mathrm{answer}}+\alpha\mathrm{CoT}_{\mathrm{max}},$
where $0\leq\alpha\leq 1$ is the hyperparameter. Empirically we identified that the best value for $\alpha$ is $0.05$ for Qwen-3B.
B.5 Detailed Results
@@ C.2 Feature importances of reasoning model
We evaluated logistic regression weights, that reflect feature importance. Our classifier achieves $0.721$, $0.717$ and $0.731$ accuracies by using thinking total entropy, length of the reasoning chain or both features combined correspondingly, thus producing a reasonable model for the evaluation of a probability of the error for question.